My Notes O Ask You (a) Calculate the absolute pressure at 101.3 kPa the bottom of a fresh water lake at a depth of 27.0 m. Assume the density of the water is 1.00 x 10 kg/m2 and the air above is at a pressure of a depth Assume the density of the water is 1.00 x 10 Pa by the water on the window of an Need Help?
Answer #1

Given

Depth of lake (h)= 27 m.

Density of water (p) = 1 × 103kg/m

Atmospheric pressure ( P, ) 101.3KP,

(a) The absolute pressure at the bottom of a fresh water lake(P).

P=P_{a}+ ho gh

101.3KP, + 1 × 10%g/m3 × 9.81m/s* × 27771

P-3.6617 × 10P

(b) Given

Diameter(D)= 33 cm

Radius(R) is

R=D/2 cm= 16.5cm= 0.165m =-

Force exerted by the water,

F=PA = P(m2)

F = 3.66 17 × 10°P, x 3.14 × (0.165111)

F = 3.13 × 104N